OpenStudy (anonymous):

Factor 5c^4 - 5c^2 - 60

OpenStudy (turingtest):

first factor out a 5. what do you have?

OpenStudy (anonymous):

c^4 - c^2 - 12

OpenStudy (turingtest):

right c^4 - c^2 - 12=0 now we call c^2=x could you factor x^2-x-12=0 ???? after we put x=c^2 back in, and see what we have.

OpenStudy (anonymous):

Still confused

OpenStudy (turingtest):

can you factor x^2-x-12=0 ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5(c^2 + 3)(c^2 - 4)

OpenStudy (turingtest):

exactly :)

OpenStudy (turingtest):

sorry I don't know why I put =0 above, ignore that bit :/

OpenStudy (anonymous):

Thks for your help :D

OpenStudy (turingtest):

welcome

OpenStudy (turingtest):

Oh sorry again, we can factor this even more by noticing that the part on the right is difference of squares:\[a^2-b^2=(a-b)(a+b)\]so\[5(c^2 + 3)(c^2 - 4)=5(c^2 + 3)(c - 2)(c+2)\]now it's completely factored.